Electric Current

Two resistances have lengths in the ratio 1:4 and radius 1:2,specific resistances in ratio 3:1. compare their resistances

vinita shinkar
13 Years agoGrade

1 Answer

Narayanarao laveti
13 Years ago

Given lengths l1=1

                         l2=4

Given radius r1=1

                        r2=2

 Given  specific resistances Ρ(rho)1=3

                                         Ρ2= 1

we know the R=Ρl/A

              R1/R2=3/1